Find the standard form of the circle x^2 + y^2 + 2x - 6y - 6 = 0
How do I solve this?
Please help.
Thank you.

Answer by stanbon(75887)

Find the standard form of the circle x^2 + y^2 + 2x - 6y - 6 = 0
--------
Complete the square on the x-terms and on the y-terms:
(x^2 +2x + 1) + (y^2 -6y + 9) = 6 +1 + 9
-------------------------------------------
Change the forms:
(x+3)^2 + (y-3)^2 = 4^2
That is the standard form.
================================
Cheers,
Stan H.